# Modified Lepage-type test statistics for the weak null hypothesis

### v1

@@ 5. $\bullet$
Table 7: Simulated actual Type I error rate and power of $\mathcal{U}(0,1)$ versus $\mathcal{U}(0,\theta)$.
$\theta$
$m$ $n$ Test 1.0 1.5 2.0 2.5 3.0
10 10 $L_{0}$ 0.0493 0.2565 0.5619 0.7444 0.8454
$L_{1}$ 0.0491 0.2794 0.5699 0.7462 0.8498
$L_{2}$ 0.0440 0.2813 0.5627 0.7524 0.8545
$L_{3}$ 0.0488 0.2417 0.5213 0.7029 0.8130
$L_{4}$ 0.0480 0.2448 0.5402 0.7318 0.8339
$L_{5}$ 0.0549 0.2556 0.5362 0.7253 0.8359
40 30 $L_{0}$ 0.0435 0.7935 0.9888 0.9993 1.0000
$L_{1}$ 0.0489 0.8122 0.9904 0.9997 1.0000
$L_{2}$ 0.0526 0.8054 0.9912 0.9994 1.0000
$L_{3}$ 0.0572 0.7874 0.9879 0.9994 1.0000
$L_{4}$ 0.0629 0.7997 0.9897 0.9994 0.9999
$L_{5}$ 0.0638 0.7993 0.9894 0.9995 1.0000
6 The empirical study
@@
The focus of this section is to explore the application and evaluation of proposed Lepage-type tests. We compare their performance to the traditional Lepage test using four illustrative datasets, seeking to identify significant changes in location, scale, or both. Figure 1 displays box-plots of the data, and Table 8 provides detailed dataset descriptions.
@@
The second dataset is taken from Hollander et al. (2014), also investigated by Marozzi (2009) and Kössler and Mukherjee (2020). This dataset originates from a biomedical experiment designed to compare peak human plasma growth hormone levels following arginine hydrochloride infusion. Group A ($m=10$), characterized by heightened cardiovascular risk and Type A personality traits, is compared with Group B ($n=11$), exhibiting resistance to coronary disease. Visual inspection (Figure 1, upper-right panel) reveals substantial inter-group differences in skewness, tail weight, location, and scale, justifying a location-scale test. The $L_{0}$, and the proposed tests, applied to this data, all yielded pp–values below 0.05 (Table 9), confirming the visual observations and aligning with prior findings (see, for example, Marozzi 2009 and Kössler and Mukherjee 2020).
@@
We investigated the sensitivity of location-scale tests to detect subtle differences in cholesterol levels related to smoking status, using Dataset 4 from the Data Story Library, a dataset also analyzed by Kössler and Mukherjee (2020). This dataset compares cholesterol levels in age-matched ex-smokers (cessation within five years) and long-term smokers (25+ years). Our research question focused on whether the Lepage-type tests could identify minor variations, even when no significant differences in mean or variance are observed, as visualized in Figure 1 (bottom-right panel). As shown in Table 9, the insignificant pp–values obtained from all competing tests indicate no difference in cholesterol levels between the groups, demonstrating the effectiveness of our proposed approaches and their alignment with traditional method.
@@
Dataset 4: Effect of smoking on cholesterol.
Smokers: 225 211 209 284 258 216 196 288 250 200 209 280
225 256 243 200 213 246 225 237 232 267 232 216
216 243 200 155 216 271 230 309 183 280 217 305
287 217 246 351 200 280 209
Ex-smokers: 250 134 300 249 213 310 175 174 328 160 188 321
213 257 292 200 271 227 238 163 263 192 242 249
242 267 243 217 267 218 217 183 228
Table 9: pp–values for the data examples.
@@
Dataset 4 0.5869 0.5879 0.5880 0.6292 0.6304 0.6304
Refer to caption Refer to caption
Refer to caption Refer to caption
Figure 1: Box-plots of the four real datasets.
@@ 7 Concluding remarks
The two-sample location-scale problem is a frequent challenge in fields such as bioinformatics and medicine. Tests that can simultaneously detect shifts in both location and scale are often more powerful than tests that only focus on location, such as the Behrens-Fisher test. This is particularly relevant in biological research, where interventions can affect both the central tendency and the variability of a dataset, requiring methods sensitive to both parameters. Given that biological data, even after preprocessing, may deviate from a normal distribution (e.g., due to outliers, heavy tails, or skewness), nonparametric approaches are often preferred. They are more robust and efficient under these conditions due to their relaxed distributional assumptions and their suitability for small to moderate sample sizes.
Lepage (1971) test statistic is a well-established nonparametric solution for the location-scale shift problem. This study introduced several new Lepage-type statistics based on different variance estimators for the Wilcoxon-Mann-Whitney ($U$) and Ansari-Bradley ($C$) statistics, which are designed for weak null hypotheses. Simulation studies across various distributions, including uniform, right-skewed, and heavy-tailed, demonstrated that the proposed tests achieve higher power in both small and large sample investigations. The application of these new statistics to four biological datasets further manifested their practical suitability. Given their enhanced power, these proposed Lepage-type statistics are recommended for practical use.
Future work is recommended in two key areas to build upon the current findings. The first involves extending the proposed methodology to include a shape parameter, which would allow for the simultaneous evaluation of location, scale, and shape. The second area of investigation is the generalization of the proposed tests for application in multi-sample comparisons.
@@
None declared.
@@
No grant received for this research.
Ethical statement
The research work has not been previously published or submitted in any form.

### v3

@@ 5.2 Power performance
Table 7: Simulated actual Type I error rate and power of $W(2,1)$ versus $W(\theta_{1},\theta_{2})$.
$\theta_{2}$
$m$ $n$ $\theta_{1}$ Test 1.0 1.2 1.5 2.0
10 10 2.0 $L_{0}$ 0.0487 0.0910 0.2353 0.6018
$L_{1}$ 0.0486 0.1007 0.2765 0.6596
$L_{2}$ 0.0486 0.1007 0.2765 0.6596
$L_{3}$ 0.0502 0.0795 0.1986 0.5346
$L_{4}$ 0.0500 0.0919 0.2475 0.6178
$L_{5}$ 0.0500 0.0919 0.2475 0.6178
2.5 $L_{0}$ 0.0612 0.1026 0.3056 0.7192
$L_{1}$ 0.0599 0.1172 0.3708 0.7895
$L_{2}$ 0.0599 0.1172 0.3708 0.7895
$L_{3}$ 0.0720 0.0909 0.2511 0.6551
$L_{4}$ 0.0646 0.1059 0.3252 0.7459
$L_{5}$ 0.0646 0.1059 0.3252 0.7459
40 30 2.0 $L_{0}$ 0.0482 0.2182 0.7771 0.9967
$L_{1}$ 0.0541 0.2341 0.7912 0.9974
$L_{2}$ 0.0544 0.2340 0.7914 0.9976
$L_{3}$ 0.0555 0.2235 0.7749 0.9964
$L_{4}$ 0.0568 0.2401 0.7881 0.9971
$L_{5}$ 0.0563 0.2396 0.7882 0.9972
2.5 $L_{0}$ 0.1078 0.2938 0.8611 0.9995
$L_{1}$ 0.1133 0.3266 0.8811 0.9996
$L_{2}$ 0.1135 0.3267 0.8815 0.9997
$L_{3}$ 0.1317 0.2998 0.8610 0.9995
$L_{4}$ 0.1374 0.3331 0.8812 0.9995
$L_{5}$ 0.1376 0.3332 0.8814 0.9996
6 Empirical applications
@@
We demonstrate the practical utility of the proposed tests using four real-world datasets from biomedical research. We compare their performance to the traditional Lepage test using four illustrative datasets, seeking to identify significant changes in location, scale, or both. Figure 1 displays box-plots of the data, and Table 8 provides detailed dataset descriptions.
@@
The second dataset is taken from Hollander et al. (2014), also investigated by Marozzi (2009) and Kössler and Mukherjee (2020). This dataset originates from a biomedical experiment designed to compare peak human plasma growth hormone levels following arginine hydrochloride infusion. Group A ($m=10$), characterized by heightened cardiovascular risk and Type A personality traits, is compared with Group B ($n=11$), exhibiting resistance to coronary disease. Visual inspection (Figure 1–upper-right panel) reveals substantial inter-group differences in skewness, tail weight, location, and scale, justifying a location-scale test. The $L_{0}$, and the proposed tests, applied to this data, all yielded pp–values below 0.05 (Table 9), confirming the visual observations and aligning with prior findings (see, for example, Marozzi 2009 and Kössler and Mukherjee 2020).
@@
We investigated the sensitivity of location-scale tests by using Dataset 4, taken from Student (1908). The analysis of the sleep duration data reveals a critical performance pattern underscoring the utility of the proposed robust test statistics. While the original Lepage test ($L_{0}$) yielded a $p$-value of $0.1257$, providing only weak evidence against the null hypothesis of distributional equality, the robust variants $L_{1}$, $L_{2}$, $L_{4}$, and $L_{5}$ produced consistently lower $p$-values ($\approx 0.0812$$0.0851$), indicating moderately stronger evidence for a location-scale alternative. This demonstrates a meaningful enhancement in sensitivity, as the robust statistics more powerfully detected the underlying pattern–where the Laevo-hyoscine group exhibited both a higher median and greater variability in sleep improvement (see, Figure 1–bottom-right panel). The lone exception, $L_{3}$, which performed similarly to $L_{0}$ ($p=0.1318$), suggests that its specific formulation of variance estimation may be less advantageous in this context. This empirical result on a classic dataset substantiates the claim that the newly developed statistics $L_{1}$, $L_{2}$, $L_{4}$, and $L_{5}$ can provide a more powerful and robust analytical tool for detecting complex distributional differences in real-world data where heteroscedasticity is present.
@@
Dataset 4: Effect of two doses on sleep hours.
Dose-1: 0.7 -1.6 -0.2 -1.2 -1 3.4 3.7 0.8 0 2
Dose-2: 1.9 0.8 1.1 0.1 -0.1 4.4 5.5 1.6 4.6 3.4
Refer to caption Refer to caption
Refer to caption Refer to caption
Figure 1: Box-plots of the four real datasets.
Table 9: pp–values for the real datasets.
@@
Dataset 4 0.1257 0.0812 0.0812 0.1318 0.0851 0.0851
@@ 7 Concluding remarks
7.1 Summary of findings
The two-sample location-scale problem represents a frequent and challenging task in fields such as biomedicine, economics, and quality control. While the classical Lepage test provides a valuable nonparametric solution, its reliance on the strong null hypothesis and underlying distributional assumptions can limit its power and robustness in practical applications, particularly with the right-skewed data prevalent in many scientific domains. This study has systematically evaluated enhanced Lepage-type test statistics that address these limitations through the integration of modern robust variance estimators.
The most significant finding from our extensive simulation study is the consistent and substantial power advantage of the proposed tests over the classical Lepage test when analyzing right-skewed data. This superior performance was robustly observed across exponential, gamma, chi-square, lognormal, and Weibull distributions. The tests incorporating both robust components ($L_{4}$, $L_{5}$) consistently demonstrated the best performance, achieving power improvements of 10-25% across various scenarios while maintaining reasonable Type I error control.
A clear performance hierarchy emerged from our results: $L_{5}\approx L_{4}\gtrsim L_{2}\approx L_{1}>L_{0}>L_{3}$, with $L_{3}$ (which modifies only the scale component) exhibiting the lowest power among the proposed tests. The equivalent performance of $L_{1}$$L_{2}$ and $L_{4}$$L_{5}$ indicates that the Fligner-Policello and Fong-Huang estimators yield virtually identical results within the Lepage statistic framework for the conditions studied. Critically, tests that robustify both location and scale components consistently outperform those modifying only the location component, which in turn surpass the classical $L_{0}$.
7.2 Limitations and methodological considerations
Our study also revealed several important limitations that warrant consideration:
Type I error inflation: Tests incorporating the new scale variance estimator ($L_{3}$, $L_{4}$, $L_{5}$) exhibited modest liberal tendencies in larger samples, with empirical sizes reaching up to 0.0594 (19% inflation). While this falls within the liberal range of the robustness criterion (see, Fagerland 2009; Morris et al. 2019), practitioners requiring strict Type I error control should use permutation-based implementations.
Distributional scope: The primary focus on right-skewed distributions, while justified by our motivating applications, limits generalizability to other distributional shapes. Our extended simulations suggest the methods perform adequately with left-skewed and symmetric with heavy-tailed distributions, but further investigation is needed.
7.3 Future research directions
Several promising directions emerge for future research:
Methodological extensions: Generalizing the methodology to accommodate a shape parameter would enable simultaneous testing of location, scale, and distributional shape. Developing multi-sample versions of these tests would greatly enhance their applicability to complex experimental designs.
Theoretical developments: Deriving finite-sample distributions, Edgeworth expansions for better approximations, and establishing optimality properties would strengthen the theoretical foundation. Investigating the asymptotic dependence structure between robustified components under alternatives represents another important direction.
Applied investigations: Exploring the performance of these statistics with other types of non-normal data, such as heavily left-skewed, multi-modal, or zero-inflated distributions, would further establish their general utility. Applications in emerging fields such as genomics, neuroimaging, and environmental statistics would demonstrate broader relevance.
7.4 Concluding statement
In conclusion, this study demonstrates that incorporating modern robust variance estimators into the Lepage framework yields tests with substantially improved performance for right-skewed data. The proposed methods offer practitioners powerful and reliable nonparametric solutions for the two-sample location-scale problem, particularly in biomedical research where right-skewed distributions are common. While theoretical and methodological challenges remain, the consistent empirical advantages shown across extensive simulations and real-data applications support their adoption in practice.
We specifically recommend the use of the $L_{2}$, $L_{4}$, or $L_{5}$ test statistics for the analysis of right-skewed data, with appropriate attention to the noted limitations and implementation considerations. Their enhanced power, achieved through robust estimation of both location and scale components, represents a meaningful advance in nonparametric methodology for practical data analysis.
We have created the LePage R package (Tsagris and Hussain, 2025), available to download from CRAN, to facilitate adoption by practitioners and researchers.
@@ 7.4 Concluding statement
The authors report there are no competing interests to declare.
@@
No specific grant was received for this research.
Data availability statement
The datasets used in the empirical study are publicly available in the cited sources.